# Resolution of Erdős Problem #728: a writeup of Aristotle’s Lean proof

### v2

@@ 1 Introduction
Figure 1: Plots showing the required inequality $v_{p}(\binom{m+k}{k})\leq\kappa_{p}(m)$; we want to find places where the red is below the blue. Both plots have $m\in[1000,2000]$ and $k=10$. The two plots are for primes $p=2$ and $p=13$, respectively. Note that these plots are smoothed by a simple moving average with window 2525, since the unsmoothed plots are dense and difficult to read. Plotting code was partially provided by ChatGPT.
@@
Acknowledgements. The author thanks forum participants in [2] for their fruitful discussion. They are, in alphabetical order, Boris Alexeev, Kevin Barreto, Thomas Bloom, Moritz Firsching, username KoishiChan, username Leeham, username old-bielefelder, Terence Tao. Specific to producing this writeup: Boris Alexeev ran Aristotle to simplify the proof; Kevin Barreto produced the original proofs; KoishiChan performed literature search; Terence Tao suggested further ideas and found extra literature.
@@ 2 Related literature and provenance
Pomerance [8] studies questions about divisors of the central binomial coefficient $\binom{2n}{n}$. One of his main theorems shows that for each fixed $k\geq 1$, the set of $n$ for which $n+k$ divides $\binom{2n}{n}$ has asymptotic density $1$ ([8, Theorem 2]). The methods in [8] exploit Kummer’s theorem and the idea that base-$p$ digit patterns behave “randomly.” Our argument is in the same vein, but differs in two key respects: (i) we treat a growing window length $k\asymp\log n$ rather than a fixed $k$, and (ii) we work with a structured divisor $\binom{m+k}{k}$ of $\binom{2m}{m}$ rather than a single linear factor $n+k$.
@@
Croot, Mousavi, and Schmidt [4] study a problem motivated by Graham’s conjecture on $\gcd\!\bigl(\binom{2n}{n},105\bigr)$. They show that for any fixed $r\geq 1$ and any $0<\varepsilon<1/(20r^{2})$, there is a threshold $p_{0}(r,\varepsilon)$ such that for any distinct primes $p_{1},\dots,p_{r}\geq p_{0}(r,\varepsilon)$ there exist infinitely many $n$ for which
@@ 3 Reduction to a valuation inequality
$\kappa_{p}(m)\ :=\ v_{p}\!\left(\binom{2m}{m}\right),\qquad W_{p}(m,k)\ :=\ v_{p}\!\Big(\prod_{i=1}^{k}(m+i)\Big),\qquad V_{p}(m,k)\ :=\ \max_{1\leq i\leq k}\ v_{p}(m+i).$
@@
$v_{p}\!\left(\binom{m+k}{k}\right)\ =\ W_{p}(m,k)-v_{p}(k!).$ (4)
@@
$v_{p}\!\left(\binom{m+k}{k}\right)\leq\kappa_{p}(m)\quad\Longleftrightarrow\quad W_{p}(m,k)\leq\kappa_{p}(m)+v_{p}(k!).$
@@
Thus, it suffices to show $W_{p}(m,k)\leq\kappa_{p}(m)+v_{p}(k!)$ for every prime $p$.
@@
The dominant contribution to $W_{p}(m,k)$ for small primes is typically $\sim k/(p-1)$, but (4) subtracts $v_{p}(k!)$, which has the same main term. This makes it easier to establish the required inequality.
@@
By Legendre’s formula, $v_{p}(k!)=\sum_{j=1}^{J}\lfloor k/p^{j}\rfloor$. This establishes the claim. ∎

### v3

@@ 1 Introduction
Figure 1: Plots showing the required inequality $\nu_{p}(\binom{m+k}{k})\leq\kappa_{p}(m)$; we want to find places where the red is below the blue. Both plots have $m\in[1000,2000]$ and $k=10$. The two plots are for primes $p=2$ and $p=13$, respectively. Note that these plots are smoothed by a simple moving average with window 2525, since the unsmoothed plots are dense and difficult to read. Plotting code was partially provided by ChatGPT.
@@
After an earlier version of this writeup was completed, Carl Pomerance has written up a note [11] extending his previous work [10] to arrive at results similar to ours. Independently, Erdős problems #729 [4] and #401 [2] were also solved by modifications of the argument in this writeup. We discuss these further developments in the second Appendix.
Acknowledgements. The author thanks forum participants in [3] for their fruitful discussion. They are, in alphabetical order, Boris Alexeev, Kevin Barreto, Thomas Bloom, Moritz Firsching, username KoishiChan, username Leeham, username old-bielefelder, Terence Tao. Specific to producing this writeup: Boris Alexeev ran Aristotle to simplify the proof; Kevin Barreto produced the original proofs; KoishiChan performed literature search; Terence Tao suggested further ideas and found extra literature. Finally, the author thanks Carl Pomerance for extending his work in light of the present developments.
@@ 2 Related literature and provenance
Pomerance [10] studies questions about divisors of the central binomial coefficient $\binom{2n}{n}$. One of his main theorems shows that for each fixed $k\geq 1$, the set of $n$ for which $n+k$ divides $\binom{2n}{n}$ has asymptotic density $1$ ([10, Theorem 2]). The methods in [10] exploit Kummer’s theorem and the idea that base-$p$ digit patterns behave “randomly.” Our argument is in the same vein, but differs in two key respects: (i) we treat a growing window length $k\asymp\log n$ rather than a fixed $k$, and (ii) we work with a structured divisor $\binom{m+k}{k}$ of $\binom{2m}{m}$ rather than a single linear factor $n+k$.
@@
Croot, Mousavi, and Schmidt [6] study a problem motivated by Graham’s conjecture on
$\gcd\!\bigl(\binom{2n}{n},105\bigr)$. They show that for any fixed $r\geq 1$ and any $0<\varepsilon<1/(20r^{2})$, there is a threshold $p_{0}(r,\varepsilon)$ such that for any distinct primes $p_{1},\dots,p_{r}\geq p_{0}(r,\varepsilon)$ there exist infinitely many $n$ for which
@@ 3 Reduction to a valuation inequality
$\kappa_{p}(m)\ :=\ \nu_{p}\!\left(\binom{2m}{m}\right),\qquad W_{p}(m,k)\ :=\ \nu_{p}\!\Big(\prod_{i=1}^{k}(m+i)\Big),\qquad V_{p}(m,k)\ :=\ \max_{1\leq i\leq k}\ \nu_{p}(m+i).$
@@
$\nu_{p}\!\left(\binom{m+k}{k}\right)\ =\ W_{p}(m,k)-\nu_{p}(k!).$ (4)
@@
$\nu_{p}\!\left(\binom{m+k}{k}\right)\leq\kappa_{p}(m)\quad\Longleftrightarrow\quad W_{p}(m,k)\leq\kappa_{p}(m)+\nu_{p}(k!).$
@@
Thus, it suffices to show $W_{p}(m,k)\leq\kappa_{p}(m)+\nu_{p}(k!)$ for every prime $p$.
@@
The dominant contribution to $W_{p}(m,k)$ for small primes is typically $\sim k/(p-1)$, but (4) subtracts $\nu_{p}(k!)$, which has the same main term. This makes it easier to establish the required inequality.
@@
By Legendre’s formula, $\nu_{p}(k!)=\sum_{j=1}^{J}\lfloor k/p^{j}\rfloor$. This establishes the claim. ∎
@@ 6 Correspondence with the Lean development
Appendix: beyond this proof
Shortly after this proof was completed, Erdős problem #729 was then solved on Jan 10, 2026, and #401 on Jan 11, 2026, again autonomously by AI. It was then subsequently noticed that these problems could be solved by modifications of the present argument as well.
Concurrently, forum participants asked Carl Pomerance, the author of [10], himself about the proof. On Jan 11, 2026, Pomerance replied that a result similar ours should be possible to obtain by modifying his argument in [10] and using a lemma which was identical to our Lemma 4. Pomerance then wrote up his extension as a note [11].
In light of these developments, it is of interest to compare our results with [11], which turn out to be very similar.
In order to unify the solutions of #728, #729, and #401 together, we must extract a general theorem out of the present argument. It is as follows.
Theorem 2.
There exist absolute constants $c_{1},c_{2}>0$ with the following properties. Consider the set $S$ of $m\in\mathbb{N}$ such that, for all $0\leq k\leq\exp(c_{1}\sqrt{\log m})$,
$\nu_{p}\left(\binom{2m}{m}\right)-\nu_{p}\left(\binom{m+k}{m}\right)\geq\left\lfloor c_{2}\frac{\log m}{\log p}\right\rfloor\cdot[p\leq 2k]$
for all primes $p$. Then $S$ has asymptotic density $1$.
Proof.
If we follow the reductions as in the writeup, to get $m\in S$ it is enough to show
$\kappa_{p}(m)-V_{p}(m,k)\geq\left\lfloor c_{2}\frac{\log m}{\log p}\right\rfloor$
for primes $p\leq 2k$.
Consider $m\in[M,2M]$ as before. Now we do not enforce $t(M)=O(\log\log M)$ but leave $t(M)$ unspecified for now. If we look at Lemma 7 we can indeed produce a gap of the form
$\left\lfloor c_{2}\frac{\log m}{\log p}\right\rfloor$
as required. The remaining thing to consider is that the union bound in Lemma 13 must still produce $|\mathrm{Bad}(M)|=o(M)$. This works out, except that we must choose $t(M)$ to be order of $\sqrt{\log M}$. Now we look back at Lemma 7 and this choice still works. ∎
To solve problem #728 [3], it is enough to show that the left-hand side in Theorem 2 is nonnegative, which is immediate from the theorem.
To solve problem #729 [4], this problem corresponds to the following. For every $c>0$, taking $k=\lfloor c\log m\rfloor$, we want to produce a threshold $p_{\min}(c)$ such that for all primes $p\geq p_{\min}$, the denominator of $(2m)!/(m!(m+k)!)$ is not divisible by $p$. By Theorem 2, we must show
$\left\lfloor c_{2}\frac{\log m}{\log p}\right\rfloor\cdot[p\leq 2k]-\nu_{p}(k!)\geq 0.$
If $p>2k$ the left-hand side is zero. So assume $p\leq 2k$. If we use $\nu_{p}(k!)\leq k/(p-1)$, it is seen that this is true for $p$ large, depending only on $c$, as required.
To solve problem #401 [2], this problem is just a more precise version of #729. The same examples work. There is an extra condition that must be satisfied which is
$\nu_{p}(m!)+\nu_{p}((m+k)!)-\nu_{p}((2m)!)\leq 2m,$
for small primes, depending on $c$. But by Legendre’s formula the left-hand side is $O(\log m)$, so this is easy to satisfy.
Finally, in [10] it was shown that
$(m+1)(m+2)\cdots(m+k)\ \Big|\ \binom{2m}{m}$
where $m$ is in a set of asymptotic density $1$, and $k$ is fixed or grows slowly with $m$. In [11], this was explicitly worked out to be $k\leq.7\log m$. We obtain a slightly weaker result, with $k\leq c\log m$ for some small constant $c>0$. The key is again the reduction to
$\left\lfloor c_{2}\frac{\log m}{\log p}\right\rfloor\cdot[p\leq 2k]-\nu_{p}(k!)\geq 0$
as in problem #729. Then we follow a similar reasoning. In [11] it was also shown that $\binom{m+k}{m}\mid\binom{2m}{m}$ for $k\leq\exp(\sqrt{\log m})$, where $m$ is in a set of asymptotic density $1$. This is slightly stronger than Theorem 2, as the constant was explicitly worked out.
As with Pomerance’s note [11], we are not sure if the bounds in Theorem 2 can be substantially improved. Nevertheless, seeing that the bound has a large jump between $p\leq 2k$ and $p>2k$, this is perhaps possible, but the author has not worked it out.
Since the subject matter of the present writeup is GPT-5.2 Pro/Aristotle’s Lean proof, it is felt that further developments will be out of scope. The reader is invited to investigate these directions further, though!